OpenStudy (anonymous):

. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4x – 8y = 48 and 11x + 3y = -105.5

OpenStudy (anonymous):

Elimination is where you are able to eliminate one of the variable terms. In this case, if you multiply the first eqn. with 3, and second eqn. with 8, you will get -24y and 24y respectively for the y terms. Then add both the eqns so these terms cancel out, and you can solve for x

OpenStudy (anonymous):

I would simplify the equations given first. 4x – 8y = 48 Divide both sides by 4 x - 2y = 12 ---(1) 11x + 3y = -105.5 ---(2) (1) x 11 - (2) 11(x - 2y) - ( 11x + 3y) = 11(12) - (-105.5) <- Solve it.